OpenStudy (anonymous):

Is there any way to find the slope of 6,8 and 6.3...?

OpenStudy (yanasidlinskiy):

Of what?

OpenStudy (yanasidlinskiy):

y=mx+b -------> Slope formula

OpenStudy (anonymous):

is there any way to find the slope of 6,8 and 6,3 or is that impossible...?

OpenStudy (anonymous):

giving medal to best answer

OpenStudy (anonymous):

its undefined (:

OpenStudy (anonymous):

your formula is y2-y1/x2-x1 for y you get -5 and for x you get 0. 5/0= undefined

OpenStudy (anonymous):

\[slope = \frac{rise}{run}=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\] in this case: \[\frac{3-8}{6-6}=\frac{-5}{0}\] Infinite slope.

OpenStudy (anonymous):

however, if you ever get an equation and you end up with something like 0/5, your answer would be 0, so be careful with that!

OpenStudy (anonymous):

I see...Thank you very much :) I was confused to be honest.

OpenStudy (anonymous):

your welcome ^.^ let me know if you need help with anything else

OpenStudy (anonymous):

Thank you again :)